Consider a hypothetical economy described by the following equations:

C = 100 + 0.8Yd


I = 1200


G = 1000


X = 1110


M = 200 + 0.25Y


T = 250 + 0.3Y


Required:


i) Compute the equilibrium level of national income. (5marks)


ii) The level of consumption, income after tax and net exports that corresponds to the equilibrium level of national/income.


1
Expert's answer
2021-06-28T06:13:01-0400

"Consumption Function is C = 100+ 0.8 Y"

 "where Y in the income in the economy."


At equilibrium level of income,

 

"AS=AD"

"Y= C+I"  

"Y= 100 + 0.8 Y + 1200"


"Y - 0.8 Y = 100 + 1200"


"0.2 Y = 1300"

"Y = 1300\/ 0.2 = 7500"


Equilibrium level of national income"= 7500"


ii) "Y = C + I + G"


"7500 = C + 1200 +1000"


"C = 7500-2200"


"level of consumption = 5300"